OpenStudy (helpmath1):

Simplify the following expression: (−3)(−2)( −2) −12 12 −18 18

jhonyy9 (jhonyy9):

bc. hope you know that minus time minus make plus so in this way -3*(-2)(-2) = -3*4 = -12 hope helped